# Fully closed maps and LUR renormability

### 06 Dec 2023 v1

@@ 1 Introduction
An equivalent LUR norm can be constructed whenever $E$ admits a separable projectional resolution of the identity. For $E=C(X)$ the simplest case is that of a compact ordinal interval (see, e.g., [2] Theorem VII.5.1). Wide classes of compacta admit a suitable transfinite sequence of retractions that corresponds to a projectional resolution of the identity. The most general of them are Valdivia compacta (see, e.g., [2] Definition VI.7.2, Theorem VI.7.6). Positive results have been obtained for some classes for which $C(X)$ does not necessarily admit a projectional resolution of the identity. Examples include $X$ scattered with $X^{(\omega_{1})}=\emptyset$ (see [11]), $X$ a compact subset of functions on a Polish space, containing countably many discontinuities (see [9]). A particular case of the latter is the Helly compact.
@@ 2 Fully closed maps and Fedorchuk compacta
In [12] Ivanov used the construction to introduce a class of compacta known as Fedorchuk compacta. We give the definition below. For some general information about inverse systems the reader is referred to [3], Chapter 2.5 or [5].
@@
In [5] II.1.6 Fedorchuk gives a list of equivalent formulations of Definition 2.1 when the spaces are regular. In the following proposition we state the three of them that will be relevant for the proof of the main result.
Proposition 2.6 ([5] II.1.6).
@@
Some authors use 2 as definition of a fully closed map (as we did in Section 1) as it is the easiest to state. If the condition of regularity is replaced by compactness, we obtain precisely Definition 1.1. Compactness would also allow for the replacement of regular in 3 with Hausdorff.
@@
Let $\pi$ be fully closed and assume that for some $f\in C(X)$ and some $\delta>0$ the set $Y_{\delta}=\left\{y\in Y\mathrel{\mathop{\ordinarycolon}}\operatorname{osc}\limits_{\pi^{-1}(y)}f>\delta\right\}$ is infinite. From compactness we can choose $y_{0}\in Y$ to be an accumulation point of $Y_{\delta}$. For $x\in\pi^{-1}(y_{0})$ choose an open neighborhood $U_{x}$ of $x$ such that $\operatorname{osc}\limits_{U_{x}}f<\delta$. The cover $\left\{U_{x}\right\}_{x\in\pi^{-1}(y_{0})}$ of the compact fiber has a finite subcover, denote it $\left\{U_{k}\right\}_{k=1}^{n}$. Then $\{y_{0}\}\cup\bigcup_{k=1}^{n}\pi^{\#}(U_{k})$ is a neighborhood of $y_{0}$ by the definition of full closedness. As $y_{0}$ is an accumulation point of $Y_{\delta}$, we can pick a point $y^{\prime}$ in $Y_{\delta}$, distinct from $y_{0}$ in this neighborhood. Then $y^{\prime}\in\pi^{\#}(U_{k})$ for some $k\in\{1,\dots,n\}$. But this implies that $\pi^{-1}(y^{\prime})\subset U_{k}$ and therefore $\operatorname{osc}\limits_{\pi^{-1}(y^{\prime})}f\leq\operatorname{osc}\limits_{U_{k}}f<\delta$ which contradicts the choice of $y^{\prime}$. To show the inverse implication assume that $\pi$ is not fully closed. By Proposition 2.6 2 there exist closed disjoint sets $F_{1},F_{2}$ in $X$ such that $f(F_{1})\cap f(F_{2})$ is infinite. By Urysohn’s lemma we can construct a continous function $f$ on $X$ satisfying $\left.f\right|_{F_{1}}=1$ and $\left.f\right|_{F_{1}}=0$. Then $\operatorname{osc}\limits_{\pi^{-1}(y)}f=1$ for all $y$ in the infinite set $f(F_{1})\cap f(F_{2})$, a contradiction. ∎
@@
A composition of fully closed mappings is not necessarily fully closed. A simple example of this fact can be found in [5]. However, the map $\pi^{2}_{0}$ sending $X_{2}$ to a point is always fully closed. The minimality condition then implies that $X_{2}$ is not metrizable. The following observation, made by Fedorchuk, gives a useful metrizability condition.
Proposition 2.8 ([5] Proposition II.3.10).
@@ 3 Preliminaries from renorming theory
Theorem 3.1 ([6], see also [17] Chapter 5).
@@ 4 Proof of the main theorem
$\tilde{f}(x)\quad=\quad\left\{\begin{aligned} \quad&f_{1}(x),\qquad&x\in\pi^{-1}\left(y_{1}\right);\\ &\vdots&\\ &f_{n}(x),\qquad&x\in\pi^{-1}\left(y_{n}\right);\\ &\tilde{g}\left(\pi(x)\right),\qquad&x\in\pi^{-1}\left(Y\setminus\bigcup_{i=1}^{n}\{y_{i}\}\right).\end{aligned}\right.$
@@ 5 Questions and remarks
It is easy to see that the lexicographic order topology on $[0,1]^{\alpha}$ coincides with that of the limit of the inverse system $\{[0,1]^{\beta},\pi^{\gamma}_{\beta},\alpha\}$, where $\pi^{\gamma}_{\beta}$ is the natural projection onto the first $\beta$ coordinates for $\beta\leq\gamma<\alpha$. The lexicographic cube $[0,1]^{\alpha}$ is thus a Fedorchuk compact of spectral height $\alpha$. The natural question that arises is the following:
@@
We will now consider an application of our main theorem. What follows is a well-known result that can be easily obtained using Theorem 3.1. Here we get it directly from Theorem 1.2.
@@
As a direct consequence we obtain the result of Deville from [1] that $C(K)$ is LUR renormable whenever $K$ is a scattered compact with $K^{(\omega)}=\emptyset$. In particular, we obtain LUR renormability of $C(K)$ for $K$ the Ciesielski-Pol compact (see [2] Examples VI.8.8). The latter is of particular significance in renorming theory as $C(K)$ is LUR renormable, but there is no linear bounded one-to-one operator from $C(K)$ into $c_{0}(\Gamma)$ for any set $\Gamma$.
@@
As above, these maps are continuous and fully closed for all $\beta<\gamma$. Moreover, for each of the projections the only nontrivial fiber is homeomorphic to the one-point compactification $\alpha\left(K^{(\beta)}\setminus K^{(\beta+1)}\right)$ of a discrete set. We can now define the inverse system $S=\left\{X_{\beta},\pi^{\beta+1}_{\beta},\gamma\right\}$. It can be shown that $S$ is a continuous inverse system and its limit $\lim\limits_{\leftarrow}S\cong K$.
In [11] Haydon and Rogers showed that $C(K)$ is LUR renormable whenever $K^{(\omega_{1})}=\emptyset$ (see also [2] Corollary VII.4.5). In this case $K$ can be reconstructed as above as the limit of a countable inverse system. This example gives rise to a more general version of Question 5.2, namely:
@@
Let $X$ be the limit of a continuous inverse system $\lim\limits_{\leftarrow}\{X_{\beta},\pi^{\beta+1}_{\beta},\gamma\}$, where $\gamma$ is a countable ordinal, $X_{0}$ is a point, the neighboring bonding mappings $\pi^{\beta+1}_{\beta}$ are fully closed, and for all $\beta<\gamma$ and all $y\in X_{\beta}$ the space $C\left((\pi^{\beta+1}_{\beta})^{-1}(y)\right)$ is LUR renormable. Does $C(X)$ admit an equivalent LUR norm?

### 24 Dec 2023 v2

@@ 1 Introduction
An equivalent LUR norm can be constructed whenever $E$ admits a separable projectional resolution of the identity. For $E=C(X)$ the simplest case is that of a compact ordinal interval (see, e.g., [2, Theorem VII.5.1]). Wide classes of compacta admit a suitable transfinite sequence of retractions that corresponds to a projectional resolution of the identity. The most general of them are Valdivia compacta (see, e.g., [2, Definition VI.7.2, Theorem VI.7.6]). Positive results have been obtained for some classes for which $C(X)$ does not necessarily admit a projectional resolution of the identity. Examples include $X$ scattered with $X^{(\omega_{1})}=\emptyset$ (see [11]), $X$ a compact subset of functions on a Polish space, containing countably many discontinuities (see [9]). A particular case of the latter is the Helly compact.
@@ 2 Fully closed maps and Fedorchuk compacta
In [12] Ivanov used the construction to introduce a class of compacta known as Fedorchuk compacta. We give the definition below. For some general information about inverse systems the reader is referred to [3, Chapter 2.5] or [5].
@@
In [5] Fedorchuk gives a list of equivalent formulations of Definition 2.1 when the spaces are regular. In the following proposition we state the three of them that will be relevant for the proof of the main result.
Proposition 2.6 ([5, II.1.6]).
@@
Some authors use (2) as definition of a fully closed map (as we did in Section 1) as it is the easiest to state. If the condition of regularity is replaced by compactness, we obtain precisely Definition 1.1. Compactness would also allow for the replacement of regular in (3) with Hausdorff.
@@
Let $\pi$ be fully closed and assume that for some $f\in C(X)$ and some $\delta>0$ the set $Y_{\delta}=\left\{y\in Y\mathrel{\mathop{\ordinarycolon}}\operatorname{osc}\limits_{\pi^{-1}(y)}f>\delta\right\}$ is infinite. From compactness we can choose $y_{0}\in Y$ to be an accumulation point of $Y_{\delta}$. For $x\in\pi^{-1}(y_{0})$ choose an open neighborhood $U_{x}$ of $x$ such that $\operatorname{osc}\limits_{U_{x}}f<\delta$. The cover $\left\{U_{x}\right\}_{x\in\pi^{-1}(y_{0})}$ of the compact fiber has a finite subcover, denote it $\left\{U_{k}\right\}_{k=1}^{n}$. Then $\{y_{0}\}\cup\bigcup_{k=1}^{n}\pi^{\#}(U_{k})$ is a neighborhood of $y_{0}$ by the definition of full closedness. As $y_{0}$ is an accumulation point of $Y_{\delta}$, we can pick a point $y^{\prime}$ in $Y_{\delta}$, distinct from $y_{0}$ in this neighborhood. Then $y^{\prime}\in\pi^{\#}(U_{k})$ for some $k\in\{1,\dots,n\}$. But this implies that $\pi^{-1}(y^{\prime})\subset U_{k}$ and therefore $\operatorname{osc}\limits_{\pi^{-1}(y^{\prime})}f\leq\operatorname{osc}\limits_{U_{k}}f<\delta$ which contradicts the choice of $y^{\prime}$. To show the inverse implication assume that $\pi$ is not fully closed. By Proposition 2.6 (2) there exist closed disjoint sets $F_{1},F_{2}$ in $X$ such that $f(F_{1})\cap f(F_{2})$ is infinite. By Urysohn’s lemma we can construct a continous function $f$ on $X$ satisfying $\left.f\right|_{F_{1}}=1$ and $\left.f\right|_{F_{2}}=0$. Then $\operatorname{osc}\limits_{\pi^{-1}(y)}f=1$ for all $y$ in the infinite set $f(F_{1})\cap f(F_{2})$, a contradiction. ∎
@@
A composition of fully closed mappings is not necessarily fully closed. A simple example of this fact can be found in [5, II.1.12]. However, the map $\pi^{2}_{0}$ sending $X_{2}$ to a point is always fully closed. The minimality condition then implies that $X_{2}$ is not metrizable. The following observation, made by Fedorchuk, gives a useful metrizability condition.
Proposition 2.8 ([5, Proposition II.3.10]).
@@ 3 Preliminaries from renorming theory
Theorem 3.1 ([6], see also [17, Chapter 5]).
@@ 4 Proof of the main theorem
$\tilde{f}(x)\quad=\quad\left\{\begin{aligned} \quad&f_{1}(x),\qquad&x\in\pi^{-1}\left(y_{1}\right);\\ &\vdots&\\ &f_{n}(x),\qquad&x\in\pi^{-1}\left(y_{n}\right);\\ &\tilde{g}\left(\pi(x)\right),\qquad&x\in\pi^{-1}\left(Y\setminus\bigcup_{i=1}^{n}\{y_{i}\}\right).\end{aligned}\right.\qed$
@@ 5 Questions and remarks
It is easy to see that for a limit ordinal $\alpha$ the lexicographic order topology on $[0,1]^{\alpha}$ coincides with that of the limit of the inverse system $\{[0,1]^{\beta},\pi^{\gamma}_{\beta},\alpha\}$, where $\pi^{\gamma}_{\beta}$ is the natural projection onto the first $\beta$ coordinates for $\beta\leq\gamma<\alpha$. The lexicographic cube $[0,1]^{\alpha}$ is thus a Fedorchuk compact of spectral height $\alpha$ (see [19]). The natural question that arises is the following:
@@
We will now consider an application of our main theorem. What follows is a well-known result that can be easily obtained using Theorem 3.1 (see e.g. [2, Corollary VII.4.6]). Here we get it directly from Theorem 1.2.
@@
As a direct consequence we obtain the result of Deville from [1] that $C(K)$ is LUR renormable whenever $K$ is a scattered compact with $K^{(\omega)}=\emptyset$. In particular, we obtain LUR renormability of $C(K)$ for $K$ the Ciesielski-Pol compact (see [2, Examples VI.8.8]). The latter is of particular significance in renorming theory as $C(K)$ is LUR renormable, but there is no linear bounded one-to-one operator from $C(K)$ into $c_{0}(\Gamma)$ for any set $\Gamma$.
@@
As above, these maps are continuous and fully closed for all $\beta<\gamma$. Moreover, for each of the projections the only nontrivial fiber is homeomorphic to the one-point compactification $\alpha\left(K^{(\beta)}\setminus K^{(\beta+1)}\right)$ of a discrete set. We can now define the inverse system $S=\left\{X_{\beta},\pi^{\alpha}_{\beta},\alpha,\beta\in\gamma\right\}$. It can be shown that $S$ is a continuous inverse system and its limit $\lim\limits_{\leftarrow}S\cong K$.
In [11] Haydon and Rogers showed that $C(K)$ is LUR renormable whenever $K^{(\omega_{1})}=\emptyset$ (see also [2, Corollary VII.4.5]). In this case $K$ can be reconstructed as above as the limit of a countable inverse system. This example gives rise to a more general version of Question 5.2, namely:
@@
Let $S=\{X_{\beta},\pi^{\alpha}_{\beta},\gamma\}$ be a continuous inverse system, where $\gamma$ is a countable ordinal, $X_{0}$ is a point, the neighboring bonding mappings $\pi^{\beta+1}_{\beta}$ are fully closed, and for any $\beta+1<\gamma$ and any $y\in X_{\beta}$ the space $C\left((\pi^{\beta+1}_{\beta})^{-1}(y)\right)$ is LUR renormable. Let $X$ be the limit $\lim\limits_{\leftarrow}S$. Does $C(X)$ admit an equivalent LUR norm?